OpenStudy (anonymous):

a university theater sold 556 tickets for a play, tickets cost $22 per adult tickets and $ 12 for senior citizen ticket, if the total receipt was 8492, how many senior citizen ticket were sold

OpenStudy (radar):

let x = number of adult tickets let y = number of senior citizens. then we know from the problem. x+y =556 and further we know the cost of each ticket so 22x + 12y=8492 Now we have two equations with two unknowns we can use substitution. x=556-y determined from first equation. now substitute in 2nd equa. 22(556-y)+12y=8492. Now solve using algebra 12232-22y+12y=8492 -10y=8492-12232 -10y=-3740 y=374 x=556-374=182 Always check (if you have time) 182 X 22 =4004 374 X 12 =4488 total 8492 Good!

OpenStudy (anonymous):

THANKS